Present: Charles, Eric, Catherine, Alex, David, Carlos, Gagik,
Jean-Marc, Matthieu, Malek, Astrid, Franck
Malek and Jean-Marc/Tagger:
All paddles have been tested with cosmics. Now will come back to paddles
which were tested "bad" during first try. This was DAQ problem though.
Analysis and decode classes are finished. Testing in progress. Next step
is comparison of results from gain match with what was done at Grenoble
for the 11 different shapes. If time permits, plan to redo everything.
Charles and Gagik/Proton Array:
Currently taking cosmic ray data for 4 towers at a time. The problem
with one tower (which was supposed to be a short or something) turned out
to be a cross-wire problem, so no disassembly was necessary. Comparison
with absolute gain calibration in ODU turns out rather badly. Geometry is
very different however (shape of blocks does matter for cosmics but not
for a LED). Monte-Carlo comparison will be coming soon. The LED pulser
board needs to feed power to 20 LED boards. It seems that the current
drawn is very high (4Amps total) which makes the VME power supply sag...
Solution is probably to plug in an external power supplu for that.
Fernando Barbosa is giving them some limited support on that. DC
monitoring has also issues, Kathy is helping to sort this out.
Catherine and Carlos/Calorimeter:
Carlos reports on his study to get find block centers. Most information
is in the Elog. Important things: position IS different depending which
LED is lit or which combination of bias for several lit LEDs. However,
position does not seem to depend on bias (good news). Also a good news,
the weighted average of the position for LED1 and LED2 and LED3 turns out
to be the same as the position of LED1+2+3 (as it should be). Conclusion
of this study by the end of the week. Catherine is writing final scripts
for getting number of photoelectrons, gains, doing linearity checks and so
on, so that it is fully automated when data taking starts. Most light
leaks have been blocked now thanks to Matthieu Jean-Marc and Alex.
David/Spectrometer:
David is working on the script to produce the HANA tree. It will be
available shortly. Gagik made a point that we should freeze (for now),
HANA's version so that everyone uses the same. Carlos will investigate
which version is most suited, most likely the last one. Also, Carlos will
write merging and Spectrometer classes in DVCS package once David has
given him the HANA output tree.
Alex and Franck/DAQ:
The change of VME CPU solved the DAQ hanging problem from last week.
Weird "empty" events noticed by Malek turns out to be a real bug for the
proton array/tagger trigger. Michel Brossard is working on that and will
fix this issue. Note that those events are only there when this trigger is
used as a real trigger. The calorimeter trigger doesn't have this issue.
Last few ARS modules will arrive next week. Our DAQ computer dvcsdaq had a
disk failure on 07/20. Had to change the disk and recover the data. Now
done. Zero-suppression mode for P-A is underway. Needs some more work,
especially getting the offsets right, and full blown test. DAQ GUI
programming underway (for now only for "newled" configuration). For now,
prestart and end run submit Elog entries with summary file. Need to make
entries in database as well. Will work on that now.
Everybody/Run Plan:
Draft run plan can be found at the following URL:
http://www.jlab.org/~sabatie/stuff/runplan.ps
Spectrometer check out will happen first. This will most likely be done
at lowest possible angle (19 degrees) where elastic scattering is still
possible (at the very limit of the momentum capabilities of the HRS left
and the a few ten Hz rates). It might actually be possible to put
protons in the proton-array in that configuration, but most likely needs
to pull back DVCS detectors a bit. Deep inelastic is also possible at low
angle with a few hundred Hz counting rates. So information will be given
to Ed Folts that Left HRS will be at 19 degrees, with negative polarity
and maximum momentum at start-up. Please verify run plan for the rest of
the settings. Colimator box issue: in or no? if we might have to use it,
knowing that we can put it in the "no collimator" setting, we should
require it + survey. Right HRS at 70 degrees. Polarity not chosen yet.
For target issues: will make a push to get at least one beer can for
hydrogen, machined cell is okay for LD2, if we can't have the beer can for
that as well. |